OpenStudy (anonymous):

13r^-4(r^-5)^5/8(r^2)^-2 solve with steps please.

OpenStudy (anonymous):

\[\frac{ 13r ^{-4}(r ^{-5})^{5} }{ 8(r ^{2})^{-2} }\]

OpenStudy (tamtoan):

use this rule : \[(a^{n})^{m} = a^{nm}\] then you can do that problem. try ?

OpenStudy (anonymous):

\[\frac{ a^n }{ a^m }=a^{n-m}\]\[a^na^m=a^{n+m}\]

OpenStudy (anonymous):

@marixoxo76 You need to read what I posted. I believe I might have misadvised you earlier I wanted to clear it up in case there was confusion.

OpenStudy (anonymous):

\[(Ax^n)(Bx^m)=(A*B)(x^{n+m})\]where A, B, n, and m are constant real numbers.